OpenStudy (ray10):

Find particular solution(integral) for the equation \[\large \frac{ d^2 x}{ dt^2 }-9x=9e ^{-12t}\] , worked out k to be \[\pm 3\] , can anyone help me for how to complete?

OpenStudy (anonymous):

So the typical solution is: \[ c_1e^{-3t}+c_2e^{3t} \]

OpenStudy (ray10):

yes that is correct :)

OpenStudy (ray10):

personally I got \[c_{1}e^{3t}+c_{2}e^{-3t}\] but it is the same :)

OpenStudy (anonymous):

So you can guess that the particular solution is\[ x_p=Ae^{-12t} \]Then integrate: \[ x_p'=-12Ae^{-12t} \]

OpenStudy (anonymous):

I mean differentiate

OpenStudy (anonymous):

Then do it a second time: \[ x''_p=144Ae^{-12t} \]

OpenStudy (ray10):

\[x''_{p}=144Ae^{-12t}\] correct

OpenStudy (ray10):

oh you have already done it :P

OpenStudy (anonymous):

YEs, now substitute \(x_p\) into the differential equation

OpenStudy (ray10):

\[144Ae^{-12t}-9Ae^{-12t}=9e^{-12t}\] is what I then get is that correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Notice that\[ 144A-9A=9 \]You can solve for \(A\).

OpenStudy (ray10):

so then once solved; \[\large A=\frac {1} {15}\] and does that make the answer \[\large A=\frac {1}{15} 9e^{-12t}\] ?

OpenStudy (ray10):

\[\large X= \frac {1}{15}9e^{-12t}\] @wio

OpenStudy (ray10):

I think that is the answer, but is there anymore to the question? :)

OpenStudy (unklerhaukus):

you had\[x_p=Ae^{-12t}\] you found A=1/15 so\[x_p=\frac1{15}e^{-12t}\]

OpenStudy (unklerhaukus):

right ? @Ray10

OpenStudy (ray10):

Oh of course!! Now I see that :D thank you @UnkleRhaukus

OpenStudy (ray10):

I have another question which I need assistance with, could you have a look at it and see if you can help me please? @UnkleRhaukus

OpenStudy (unklerhaukus):

i sure can

OpenStudy (ray10):

I shall post a new question of it :)